At 6.4 percent interest, how many years does it take to double your money? (round to 2 decimal places)
Answer #1

ANswer : 11.17 Years

Year to Take the Double the Money At the rate of 6.4 % per annum

assume = Current Money is $ 100

It Must be Double = 2*100 = 200

At 6.4% interest rate

Future Value = I*(1+i)^n

200 = 100*(1.064)^n

n= 11.174

200= 100*(1.064)^11.174

Answer : 11.174Years

Note: IF we follow Rule 72 ( Short cut method to Find out the duration)

Years to Take the double the Money = 72/ i = 72/6.4 = 11.25 Years